[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Tue Jul 25, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Wed Jul 26 03:07:05 UTC 2017


HSU Shift Log for Tue Jul 25 2017 23:06:59 GMT-0400 (EDT)

Shift: R

Forecaster: SR Stewart

Technical Issues Summary:

ATCF issues: None (see separate JTWC issues below).

N-AWIPS issues: None.

AWIPS-2/NCP issues: None.

Other Techincal issues: (1) JTWC called at 2200Z and said they were still  
having problems transferring ATCF files for all three EPAC storms; Monica  
and Matt Sardi were notified, and technical problems were resolved. (2) the  
FSSE model arrived, but it was too late for the H. Hilary and H. Irwin 03Z  
advisories.

Miscellaneous issues: Andy Latto (ASL) from TAFB worked HSM training shift,  
and ATCF & forecast training was provided by SRS and JPC; requiring only  
with limited assistance, ASL prepared forecast/advisory package #35 on TD  
Greg.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170726/bc177aa6/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list